A 110mm diameter cylinder contains 100cm3 of water at 60°C. A 50kg piston sits on top of the water. If heat is added until the temperature is 200°C, find the work done.
1
Expert's answer
2019-06-20T11:16:16-0400

The work is given by formula

"A=p(V_2 - V_1) (1)"

where p is the pressure

We can write for force balance in cylinder


"mg=pS \u2013 p_{atm} S (2)"

where m=50kg, g=10 N/kg, patm=105 Pa

Area of the piston is equal to

"S=\frac { \pi d^2} {4} (3)"

We put (3) in (2) and get: p=151600 Pa

Gay-Lussac's law is given by formula

"\frac {V_1} {V_2} = \frac {T_1} {T_2} (4)"

where V1=0.0001 m3, T1=333 K, T2=473 K

Using (4) we get: V2=0.142 m3

According to formula (1), we calculate the work:

A=21.5 kJ


Answer:

21.5 kJ